# Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v1

@@ 5 Related Work
5.0.1 Prediction with Rejection.
Selective classification falls into the broader problem setting of prediction with rejection. In the case of SC, misclassifications are to be rejected [14]. The baseline approach is to use the MSP [30, 22] and there have been a number of proposed training [57, 33, 89] and architectural [23, 11] enhancements, however, recently the effectiveness of some of these enhancements has been called into question [15]. In this work, we focus on only the MSP. Another scenario involving prediction with rejection is out-of-distribution (OOD) detection [84, 30], where data from outside of the training distribution are to be rejected. There is a plethora of varied research in this field [82, 68, 49, 88, 50, 34, 77, 29] extending to semantic segmentation [26, 46, 2, 24, 18]. Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data [36, 81, 38, 5]. Notably, Deep Ensembles [43] have arisen as a reliable method to improve performance in all three scenarios [38, 83, 45]. We believe, given the results of this work, that extending the investigation of LS (and other training enhancements) to other scenarios involving prediction with rejection is an important avenue of future work.
5.0.2 Mixup.
Mixup [87] and its variants [86, 17, 62, 54, 3] are a set of regularisation techniques that involve interpolating between random pairs of samples at training time, modifying both inputs and targets. They are commonly found in many ImageNet training recipes [73, 52, 51] and often used in conjunction with label smoothing [80, 74, 53]. We remark that research into how Mixup and its variants affect SC is a particularly salient avenue of future work as it similarly “softens” the classification training targets.
5.0.3 Label Smoothing.
Beyond prediction with rejection, label smoothing [70] is well investigated. It has been shown to improve model calibration [59, 58, 47, 10] as well as accuracy when training under label noise [56]. Knowledge distillation [31], where soft labels are provided by a teacher network, is also commonly linked and combined with LS [59, 21, 65, 6, 85] due to their similarity. Interestingly, in a similar vein to our work, pre-training using LS has been shown to harm transfer learning [40].
@@ 6 Concluding Remarks
In this work, we investigate the effect that label smoothing has on a model’s ability to perform selective classification. Our experiments across a range of architectures and tasks reveal that the use of label smoothing leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves classification performance. By analysing the logit-level gradients of the label smoothing loss, we are able to explain this behaviour – label smoothing increasingly regularises the max logit as the true probability of error decreases, exacerbating overconfidence and underconfidence. We then investigate logit normalisation as a potential method to improve the degraded selective performance of label-smoothing-trained models. We find it to be highly effective, as it, in essence, reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher. We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation.
6.0.1 Acknowledgments.
This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-[AD011011970R3]). GX’s PhD is funded jointly by Arm and the EPSRC.
Supplementary Material
Appendix 0.A Additional Results
@@
  • Fig. 11 shows ViT-S-16 evaluated on the combination of ImageNet and ImageNet-Sketch. The behaviour mirrors that of ResNet-50 in the main paper.
  • Fig. 12 shows the effectiveness of logit normalisation on DeepLabV3+ (ResNet-101) on Cityscapes.
  • Figs. 13 and 14 shows the distribution of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ViT-S-16 and DeepLabV3+. We see that similarly to ResNet-50, for LS, the distribution of $v_{\text{max}}$ is higher for errors ✗. Although it is less obvious, it is clear for both ViT-S-16 and DeepLabV3+ that for higher $\pi_{\text{max}}$ the standard deviations overlap much less than for CE.
@10% coverage of ImageNet + Sketch
ImageNet
Sketch
CE
#samples
9528
560
#errors
88
43
error rate
0.9
7.7
LS $\alpha=0.1$
#samples
9332
756
#errors
130
83
error rate
1.4
11.0
LS $\alpha=0.2$
#samples
9117
971
#errors
177
164
error rate
1.9
16.9
LS $\alpha=0.3$
#samples
8966
1122
#errors
208
195
error rate
2.3
17.4
Figure 11: ViT-S-16 evaluated on the combination of ImageNet and ImageNet-Sketch. The results mirror ResNet-50 from the main paper. LS leads to increased overconfidence when epistemic uncertainty is introduced, as shown through the increase in the failure rate amongst high-confidence samples.
Figure 12: RC curve showing the effect of logit normalisation for DeepLabV3+ (ResNet-101). The behaviour is similar to the results in the main paper. Logit normalisation is effective in improving the performance of the LS-trained model, however, makes little difference to the CE-trained model.
Figure 13: The max logit $v_{\text{max}}$ as a function of the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately for ViT-S-16. Similarly to the main paper, the distribution is higher for errors  for the LS-trained model. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
@@
0.A.1 Small-Scale CIFAR Experiments
We also include experimental results on small-scale $32\times 32$ CIFAR-100 [41]. We train a DenseNet-BC [32] ($k=12,L=100$) to show further generality over model architecture families. Figs. 15, 16 and 17 show results that mirror those found in the main paper for ResNet-50 on ImageNet, although we note that LS does not improve top-1 error rate in this case.
@@ 0.A.1 Small-Scale CIFAR Experiments
Appendix 0.B Reproducibility
Alongside this document, we provide a code demo to train two ResNet-20 [27] on CIFAR-10 [41] with cross-entropy and label smoothing and compare the corresponding Risk-Coverage curves. We recall that all the code used in this project will be made available shortly to ensure exact reproducibility. We also plan to release our most important models on Hugging Face.
Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix 0.A. All our models are trained with PyTorch, and we use the native implementation from the CrossEntropyLoss for label smoothing.
0.B.1 Image Classification
@@
For the DenseNet trained on CIFAR-100, we extract a subset of CIFAR’s training set containing 5000 images to create a proper validation set. We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov [60, 69], a momentum of 0.9, and $1\times 10^{-4}$ weight decay. We divide the learning rate by ten after 150 and 225 epochs. We use standard augmentations. We apply random crop with a four-pixel padding as well as random horizontal flip. We do not perform model selection and keep the last checkpoint.
ViT-16-S – ImageNet.
For our ViT-16-S trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW [55] with the $\beta$s equal to 0.9 and 0.999. We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate. The models are trained with non-adaptive sharpness-aware minimisation (SAM) [16, 8] with $\rho=0.2$. We use a dropout [66] rate of 0.1 but no attention dropout. We transform the training images with a standard random resized crop to 224$\times$224 pixels using bicubic interpolation and a random horizontal flip. For evaluation, we center-crop the images to this resolution. For ImageNet, we do not perform model selection and keep the last checkpoint. However, we randomly extract a validation set of 50000 images from the training set to perform logit normalisation.
@@
Our ResNet-50 is trained on 8 V100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images. After five epochs of linear warmup, we use a cosine annealing scheduler starting with a learning rate of 0.4 with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We use the same transformation of the images as for the ViT and select the last checkpoint for inference. We use the same evaluation set as for the ViT for logit normalisation.
0.B.2 Semantic Segmentation
@@
We train a Deeplabv3+ on CityScapes [12] with a ResNet-101 [27] backbone pre-trained on ImageNet [63]. We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy [48] with a power of 0.9. The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$. We take a batch size of 12 images and train for 40000 steps. During training, we randomly crop the input images and targets to squares with 768-pixel-long sides. We apply random horizontal flip and colour-jitter with the classical parameters: brightness, contrast, and saturation levels of 0.5. For testing, we use the images at their original resolution and do not perform any test time augmentations. For the RC curves, we randomly sample 5000 predictions per image, extracted prior to the final interpolation to reduce correlations between the predictions. We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
Appendix 0.C Result and Proof
@@
Appendix 0.D Discussions
0.D.1 On the Assumption that Logits are Positive

### v2

@@ 6 Related work
Prediction with rejection. Selective classification falls into the broader problem setting of prediction with rejection. In the case of SC, misclassifications are to be rejected (El-Yaniv & Wiener, 2010). The baseline approach is to use the MSP (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017) and there have been a number of proposed training (Moon et al., 2020; Huang et al., 2020; Ziyin et al., 2019; Zhu et al., 2024) and architectural (Geifman & El-Yaniv, 2019; Corbière et al., 2019) enhancements, however, recently the effectiveness of some of these enhancements has been called into question (Feng et al., 2023). Another scenario is out-of-distribution (OOD) detection (Yang et al., 2021; Hendrycks & Gimpel, 2017), where data from outside of the training distribution are to be rejected. There is a plethora of research in this field (Xia & Bouganis, 2022a; Sun et al., 2021; Liu et al., 2020; Zhang et al., 2023; Liu et al., 2023; Wang et al., 2022; Hendrycks et al., 2022). Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data (Jaeger et al., 2023; Xia & Bouganis, 2022b; Kim et al., 2023). Notably, Deep Ensembles (Lakshminarayanan et al., 2020) have arisen as a reliable method to improve performance in all three scenarios (Kim et al., 2023; Xia & Bouganis, 2023; Laurent et al., 2023; 2024). We believe, given the results of this work, that extending the investigation of LS (and other training enhancements) to other scenarios involving prediction with rejection is an important avenue of future work.
Mixup. Mixup (Zhang et al., 2018) and its variants (Yun et al., 2019; Franchi et al., 2021; Pinto et al., 2022; Liu et al., 2022d; Bouniot et al., 2023) are a set of regularisation techniques that involve interpolating between random pairs of samples at training time, modifying both inputs and targets. They are commonly found in many ImageNet training recipes (Tan & Le, 2021; Liu et al., 2021; 2022b) and are often used in conjunction with label smoothing (Wightman et al., 2021; Touvron et al., 2021; Liu et al., 2022c). Research into how Mixup affects SC is a particularly salient avenue of future work as (Zhu et al., 2024) also observe empirically that it can have a negative impact.
Label smoothing. Beyond prediction with rejection, LS is well-explored across various contexts. It has been shown to improve model calibration (Müller et al., 2019; Chun et al., 2020; Mukhoti et al., 2020; Liu et al., 2022a) as well as accuracy when training under label noise (Lukasik et al., 2020). Knowledge distillation (Hinton et al., 2015), where soft labels are provided by a teacher network, is also commonly linked and combined with LS (Müller et al., 2019; Gao et al., 2020; Yuan et al., 2020; Shen et al., 2021; Chandrasegaran et al., 2022) due to their similarity. Interestingly, in a similar vein to our work, pre-training using LS has been shown to harm transfer learning (Kornblith et al., 2021).
@@ 7 Concluding remarks
In this work, we elucidate the effect of label smoothing (LS) on selective classification (SC). Our experiments across various tasks and architectures show that LS leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves accuracy. By analysing the logit-level gradients of the LS loss, we provide an explanation for this previously not understood behaviour – LS regularises/suppresses the max logit more(less) the more likely a prediction will be correct(wrong), degrading a model’s uncertainty ranking of correct ✓ vs incorrect ✗ predictions. We then investigate post-hoc logit normalisation as a method to improve the degraded SC performance caused by LS. We find it to be highly effective and shed light on why – it reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher. We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation.
Reproducibility statement
To ensure transparency, we use publicly available datasets: CIFAR-100, ImageNet, and Cityscapes. Our detailed experimental methods are outlined in Appendix B. Finally, the proof supporting our explanation of the effectiveness of logit normalisation is provided in Appendix D.
To help replicate our work, we will soon share our source code on GitHub, including configuration files and the training code. We also plan to release our most important models on Hugging Face.
Ethics
Our primary objective is to contribute to enhancing the reliability of machine learning methods, with a particular focus on raising awareness about the negative effects of label-smoothing on selective classification. While our work aims to improve the robustness of ML systems, we acknowledge the potential risk that these advancements could be misapplied in harmful ways.
Appendix A Additional results
A.1 Illustrative examples of ImageNet overconfidence
Fig. 9 shows a few examples of overconfident misclassifications on the ImageNet evaluation data by our LS-trained ResNet-50.
Figure 9: Illustrative examples of overconfident errors performed by our LS-trained ResNet-50 on evaluation data. Even though $P_{\text{error}}$ is high in all cases (e.g. due to multiple possible labels), the model predicts low (ranking) uncertainty. Note that even though the CE model is wrong as well, it has assigned higher ranking uncertainties, reflecting its superior SC ability shown in Fig. 3.
A.2 Complete results for ViT and DeepLabV3+
@@
  • Fig. 10 and Tab. 1 show full ResNet-50 and ViT-S-16 results on ImageNet + ImageNet-Sketch. We see that the behaviour of ViT-S-16 is similar to ResNet-50 (increasing numbers of confident errors from sketch as $\alpha$ increases), but less pronounced.
  • Figs. 11 and 12 shows the distribution of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ViT-S-16 and DeepLabV3+. We see that similarly to ResNet-50, for LS, the distribution of $v_{\text{max}}$ is higher for errors ✗. Although it is less obvious, it is clear for both ViT-S-16 and DeepLabV3+ that for higher $\pi_{\text{max}}$ the standard deviations overlap much less than for CE.
  • Fig. 13 shows the effectiveness of logit normalisation for ViT-S-16 on ImageNet and DeepLabV3+ (ResNet-101) on Cityscapes. We also provide 2 additional segmentation figures in Appendix F.
Figure 10: RC curves when evaluating on the combination of ImageNet and ImageNet-sketch. We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced. This shows that LS leads to increasing overconfidence on less-well-fit data.
Table 1: Statistics @10% coverage of the combined evaluation set. As the level of LS $\alpha$ increases, the number of errors increases, especially the number of errors from ImageNet-Sketch.
ResNet-50 @10% coverage of ImageNet + Sketch
ImageNet Sketch
CE #samples 8994 1094
#errors 93 80
error rate 1.0 7.3
LS $\alpha=0.1$samples 8561 1527
#errors 187 211
error rate 2.2 13.8
LS $\alpha=0.2$samples 8171 1917
#errors 244 365
error rate 3.0 19.0
LS $\alpha=0.3$samples 7915 2173
#errors 297 541
error rate 3.8 24.9
ViT-S-16 @10% coverage of ImageNet + Sketch
ImageNet Sketch
CE #samples 9528 560
#errors 88 43
error rate 0.9 7.7
LS $\alpha=0.1$samples 9332 756
#errors 130 83
error rate 1.4 11.0
LS $\alpha=0.2$samples 9117 971
#errors 177 164
error rate 1.9 16.9
LS $\alpha=0.3$samples 8966 1122
#errors 208 195
error rate 2.3 17.4
Figure 11: Distribution of the max logit $v_{\text{max}}$ given MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately for ViT-S-16. Similarly to the main paper, the distribution of errors  is higher than correct predictions  for the LS-trained model. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
@@
Figure 13: RC curve showing the effect of logit normalisation for ViT-S-16 and DeepLabV3+ (ResNet-101). The behaviour is similar to the results in the main paper. Logit normalisation is effective in improving the performance of the LS-trained models, bringing them close to the CE models. However, logit normalisation makes little difference to the CE-trained model.
A.3 Small-scale CIFAR experiments
We also include experimental results on small-scale $32\times 32$ CIFAR-100 (Krizhevsky, 2009). We train a DenseNet-BC (Huang et al., 2017) ($k=12,L=100$) to show further generality over model architecture families. Figs. 14, 15 and 16 show results that mirror those found in the main paper for ResNet-50 on ImageNet, although we note that LS does not improve top-1 error rate in this case.
@@ A.3 Small-scale CIFAR experiments
Appendix B Reproducibility
We recall that all the code used in this project will be made available soon to ensure exact reproducibility. We also plan to release our most important models on Hugging Face.
Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix A. All our models are trained with PyTorch, and we use the native implementation from the CrossEntropyLoss for label smoothing.
B.1 Image classification
@@
For the DenseNet trained on CIFAR-100, we extract a subset of CIFAR’s training set containing 5000 images to create a proper validation set. We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov (Nesterov, 1983; Sutskever et al., 2013), a momentum of 0.9, and $1\times 10^{-4}$ weight decay. We divide the learning rate by ten after 150 and 225 epochs. We use standard augmentations. We apply random crop with a four-pixel padding as well as random horizontal flip. We do not perform model selection and keep the last checkpoint.
ViT-S-16 – ImageNet.
For our ViT-S-16 trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW (Loshchilov & Hutter, 2019) with the $\beta$s equal to 0.9 and 0.999. We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate. The models are trained with non-adaptive sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022) with $\rho=0.2$. We use a dropout (Srivastava et al., 2014) rate of 0.1 but no attention dropout. We transform the training images with a standard random resized crop to 224$\times$224 pixels using bicubic interpolation and a random horizontal flip. For evaluation, we center-crop the images to this resolution. For ImageNet, we do not perform model selection and keep the last checkpoint. However, we randomly extract a validation set of 50000 images from the training set to perform logit normalisation.
@@
Our ResNet-50 is trained on 8 V100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images. After five epochs of linear warmup, we use a cosine annealing scheduler starting with a learning rate of 0.4 with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We use the same transformation of the images as for the ViT and select the last checkpoint for inference. We use the same validation set as for the ViT for logit normalisation.
B.2 Semantic segmentation
@@
We train a Deeplabv3+ on CityScapes (Cordts et al., 2016) with a ResNet-101 (He et al., 2016) backbone pre-trained on ImageNet (Russakovsky et al., 2015). We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy (Liu et al., 2015) with a power of 0.9. The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$. We take a batch size of 12 images and train for 40000 steps. During training, we randomly crop the input images and targets to squares with 768-pixel-long sides. We apply random horizontal flip and colour-jitter with the classical parameters: brightness, contrast, and saturation levels of 0.5. For testing, we use the images at their original resolution and do not perform any test time augmentations. For the RC curves, we randomly sample 5000 predictions per image, extracted prior to the final interpolation to reduce correlations between the predictions. We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
B.3 Logit Normalisation
[…]
Appendix C Analysis of empirical loss (one-hot labels)
We can perform a similar analysis as in Sec. 3 using the empirical loss (Eqs. 2 and 6) rather than $\mathcal{L}^{\text{true}}$ as in the main paper. We leave this to the Appendix as the conclusions are similar to analysing $\mathcal{L}^{\text{true}}$, however, the presence of one-hot targets makes it less convenient to reason about the probability of error $P_{\text{error}}$ and how well fit the model is to the true conditional distribution $P_{\text{data}}(y|\boldsymbol{x})$ for a given data sample. Taking the gradients of Eqs. 2 and 6 as in Eq. 13,
$\frac{\partial\mathcal{L}_{\text{CE}}}{\partial v_{k}}=-\left[\delta_{y\omega_{k}}\!\!-\pi_{k}\right],\quad\frac{\partial\mathcal{L}_{\text{LS}}}{\partial v_{k}}=-\biggl{[}\Bigl{[}\,\underbrace{(1-\alpha)\delta_{y\omega_{k}}}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}+\!\!\!\!\underbrace{\alpha/K}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!\Bigr{]}\!-\pi_{k}\biggr{]},$ (19)
which in turn gives the regularisation gradient,
$\frac{\partial\mathcal{L}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}-\mathcal{L}_{\text{CE}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}}{\partial v_{k}}=\alpha\left[\delta_{y\omega_{k}}-1/K\right]~{},$ (20)
which is once again independent of the model output $\boldsymbol{\pi}$. In this case, as the label $y$ has already been sampled, the model is either right  or wrong , giving two different regularisation gradients for the max logit $v_{\text{max}}$,
[…]
We can see that the max logit is more strongly regularised/suppressed during training when the prediction is correct, which aligns with the analysis of $\mathcal{L}^{\text{true}}$ and $P_{\text{error}}$ in the main paper. When the model is correct during training, the max logit is suppressed, but when it is incorrect the max logit is not suppressed. Thus the uncertainty ranking of correct ✓ vs incorrect ✗ data samples is degraded, harming SC. This leads us to the same conclusions as in Sec. 4 and also aligns with the behaviour in Fig. 5 (right) where $v_{\text{max}}$ is higher for errors given the value of MSP.
Appendix D Result and proof
@@
Appendix E Discussions
E.1 $U$ other than MSP
In the main body of the paper we focus solely on MSP as our uncertainty score $U$. Fig. 17 shows how LS affects the SC behaviour of MSP (top) compared two other softmax scores: DOCTOR (Granese et al., 2021) ($U=-||\boldsymbol{\pi}||_{2}$) (middle) and entropy ($U=H(\boldsymbol{\pi})=-\sum_{k}\pi_{k}\log\pi_{k}$) (bottom). We see that the behaviour is very similar across all three softmax scores, with increasing the level of LS $\alpha$ degrading all the scores. As MSP is the (marginally) best performing and the most commonly used, we thus choose to focus on it for the main paper.
Fig. 18 shows the SC performance of Energy ($U=-\log\sum_{k}\exp v_{k}$) (Liu et al., 2020), a popular OOD detection score, compared to MSP with and without LS. We see that Energy performs much worse than MSP at SC. This aligns with a large body of existing work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Kim et al., 2023; Yang et al., 2024; Zhu et al., 2024) that empirically finds that uncertainty scores designed for OOD detection perform poorly at detecting misclassifications and are thus not suitable for SC. Thus we choose not to investigate any OOD detection scores in the main paper. We remark that LS also has a strong negative effect on the SC performance of Energy. This is unsurprising as Energy is dominated by the max logit. An important future direction would be to investigate the effect of LS on various OOD detection scores on the task of OOD detection.
Finally, we also choose to omit various training/architecture-based approaches such as (Ziyin et al., 2019; Geifman & El-Yaniv, 2019; Moon et al., 2020; Zhu et al., 2024), as we aim to focus our investigation solely on understanding the training effects of LS. We note that LS (potentially combined with post-hoc logit normalisation) is considerably simpler to implement than the aforementioned training-based approaches, and has been shown to be stable in many more use-cases and so may be more likely to be chosen by a practitioner. Besides, recent work (Feng et al., 2023) has suggested that MSP applied to the methods in (Ziyin et al., 2019; Geifman & El-Yaniv, 2019) actually performs better than the $U$s proposed in them (reject logit/selection head).
Figure 17: The effect of LS on different softmax scores: MSP (top), DOCTOR (middle), Entropy (bottom). We see that the behaviour is very similar, with MSP being the best performer.
Figure 18: SC performance of MSP and Energy (OOD detection score). Energy significantly underperforms MSP. This behaviour is in line with existing work that shows that uncertainty scores designed for OOD detection are not suitable for SC.
E.2 Aleatoric and epistemic uncertainty
We recognise that the conceptual decomposition of predictive uncertainty into aleatoric and epistemic uncertainty (Gal, 2016; Hüllermeier & Waegeman, 2021; Kirsch, 2024) can be applied to the discussion throughout this paper, and that some readers may be confused as to why we do not use these terms. We choose to use simpler, more direct language as we believe it more efficiently conveys our discussion, reduces the number of concepts to introduce and also reduces any potential confusion.
E.3 On the assumption that logits are positive